# On the relationship between Koopman operator approximations and neural ordinary differential equations for data-driven time-evolution predictions

### v1

@@ III Numerical Methods
where $\mathbf{x}_{i}\equiv\mathbf{x}(t_{i})$, $\tau=t_{i+1}-t_{i}$ is the sampling interval, and $m$ is the number of data pairs. We note that for the continuous time formulations considered below, such as neural ODEs, it is not necessary to have a fixed time interval between datapoints in the dataset, but we will take it to be fixed here for consistency between methods. We will additionally be interested in fitting models using neural networks as function approximators, and those which preserve the Markovian structure assumed in the dynamical systems formulation above – that future time evolution is determined only from the present states. This requirement excludes several popular deep learning based time series modeling approaches, such as reservoir computing, LSTMs, and transformers.
@@ III.1 Neural ODEs
Perhaps the most straightforward way to construct a dynamics model from data is to pose a functional form for the flow map in Eq. 2, as
$\mathbf{x}_{i+1}=\mathbf{F}_{\tau}(\mathbf{x}_{i};\theta_{\mathbf{F}})$ (9)
where the parameters $\theta_{\mathbf{F}}$ are chosen to minimize the prediction loss
$J(\theta_{\mathbf{F}})=\sum_{i=1}^{m}\|\mathbf{y}_{i}-\mathbf{F}_{\tau}(\mathbf{x}_{i};\theta_{\mathbf{F}})\|_{2}^{2}\,.$ (10)
For example, $\mathbf{F}_{\tau}$ can be represented as a simple feedforward neural network and the objective minimized by stochastic gradient descent. Alternatively, $\mathbf{F}_{\tau}$ could be assumed to be linear in a predefined set of features $\psi_{j}(\mathbf{x})$, such as monomials, radial basis functions, or a Fourier basis; and then the parameters can be optimized by solving a linear least squares problem.
In many cases, we would prefer a continuous time model, which requires an approximation of the ODE vector field $\mathbf{f}(\mathbf{x})$ in Eq. 1. When $\mathbf{f}$ is approximated by a neural network $\mathbf{f}(\mathbf{x};\theta_{\mathbf{f}})$, the resulting model is called a neural ODE [50]. The parameters of a neural ODE are optimized to minimize the objective
@@
The objective in Eq. 11 is typically optimized by stochastic gradient descent, where the gradients with respect to the parameters $\theta_{\mathbf{f}}$ are determined by either solving an adjoint equation backward in time from $t_{i}+\tau$ to $t_{i}$ to obtain these sensitivities or by using modern automatic differentiation tools to backpropagate these gradients through the operations of a numerical ODE solver [50]. Once the neural ODE model is trained, it can be implemented to predict the time evolution of a system by simply integrating along the learned vector field using a numerical ODE solver from a given initial condition.
III.2 EDMD
Extended dynamic mode decomposition (EDMD) [21] is a method for constructing a matrix approximation of the Koopman operator from data by considering the operator acting on elements of a function subspace spanned by a predefined set of dictionary functions. Consider the dictionary
$\mathbb{D}=\{\psi_{j}:\mathbb{R}^{n}\to\mathbb{R}\}_{j=1}^{k}$ (13)
whose elements span a subspace $\mathcal{O}_{\mathbb{D}}\subset\mathcal{O}$. Any observable $g_{\mathbf{c}}\in\mathcal{O}_{\mathbb{D}}$ can be represented as a linear combination of the dictionary elements
$g_{\mathbf{c}}(\mathbf{x})=\sum_{j=1}^{k}\mathbf{c}_{j}\psi_{j}(\mathbf{x})=\mathbf{c}^{\mathrm{T}}\Psi(\mathbf{x})$ (14)
where $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{k}$ is a vector-valued function containing the elements of $\mathbb{D}$, and $\mathbf{c}\in\mathbb{R}^{k}$ is a vector of projection coefficients. A time-evolved observable $\mathcal{K}_{\tau}g_{\mathbf{c}}$ can also be expressed in terms of the dictionary elements as
$\mathcal{K}_{\tau}g_{\mathbf{c}}(\mathbf{x})=\mathbf{c}^{\mathrm{T}}\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x})+r_{\mathbf{c}}(\mathbf{x})$ (15)
where $\mathbf{K}_{\tau}\in\mathbb{R}^{k\times k}$ is a matrix which approximates the action of the Koopman operator $\mathcal{K}_{\tau}$ by updating the projection coefficients $\mathbf{c}$ and $r_{\mathbf{c}}\in\mathcal{O}$ is a residual, which we will seek to minimize by an appropriate choice of $\mathbf{K}_{\tau}$. If $\mathcal{O}_{\mathbb{D}}$ is Koopman invariant, then the operator can be represented exactly as a matrix and the residual, $r_{\mathbf{c}}(\mathbf{x})=0$. Note that the time-evolved observable can also be represented, using Eq. 3, as
$\mathcal{K}_{\tau}g_{\mathbf{c}}(\mathbf{x})=g_{\mathbf{c}}\circ\mathbf{F}_{\tau}(\mathbf{x})=\mathbf{c}^{\mathrm{T}}\Psi\circ\mathbf{F}_{\tau}(\mathbf{x}).$ (16)
Equating Eq. 15 and Eq. 16, the residual $r_{\mathbf{c}}(\mathbf{x})$ can be written as
$r_{\mathbf{c}}(\mathbf{x})=\mathbf{c}^{\mathrm{T}}\big{(}\Psi(\mathbf{F}_{\tau}(\mathbf{x}))-\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x})\big{)}\,.$ (17)
Since we would like to minimize the residual for arbitrary observables in $\mathcal{O}_{\mathbb{D}}$, we can drop the dependence on $\mathbf{c}$ and seek to minimize the objective
@@ III.2 EDMD
where $\mathbf{y}_{i}=\mathbf{x}_{i+1}=\mathbf{F}_{\tau}(\mathbf{x}_{i})$. If the lifted measurement data is organized into the lifted data matrices $\Psi_{X},\Psi_{Y}\in\mathbb{R}^{k\times m}$ as
$\Psi_{X}=\begin{bmatrix}\Psi(\mathbf{x}_{1})&,~{}\cdots~{},&\Psi(\mathbf{x}_{m})\end{bmatrix}\qquad,\qquad\Psi_{Y}=\begin{bmatrix}\Psi(\mathbf{y}_{1})&,~{}\cdots~{},&\Psi(\mathbf{y}_{m})\end{bmatrix}$
then $J(\mathbf{K}_{\tau})$ can be expressed succinctly as
$J(\mathbf{K}_{\tau})=\|\Psi_{Y}-\mathbf{K}_{\tau}^{\mathrm{T}}\Psi_{X}\|_{F}^{2}\,$ (19)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. The solution of this least-squares problem for $\mathbf{K}$ is given by
$\mathbf{K}_{\tau}^{\mathrm{T}}=\Psi_{Y}\Psi_{X}^{\mathrm{T}}\left(\Psi_{X}\Psi_{X}^{\mathrm{T}}\right)^{\dagger}$ (20)
where $(\cdot)^{\dagger}$ denotes the Moore-Penrose pseudoinverse [21]. We further detail how such a matrix approximation can be used for predicting the time evolution of the states below in Sec. III.4.
III.3 EDMD with dictionary learning
A challenge of the EDMD approach is finding an appropriate dictionary $\mathbb{D}$ on which to compute the operator $\mathbf{K}_{t}$, as the as the accuracy and convergence of the approximation depend heavily on this choice. Ideally, the subspace spanned by the dictionary should be Koopman invariant so that the residual can be driven to zero. To meet this requirement, it is necessary that the dictionary elements all lie in a space spanned by Koopman eigenfunctions [12, 16]. Additionally, we would like a dictionary from which the states can be recovered easily, and preferably linearly, for use in prediction applications. However, finding a dictionary that satisfies these properties is nontrivial, as the appropriate dictionary depends heavily on the underlying dynamical system and many common choices of dictionary do not scale well with the state dimension, making them challenging to apply to systems of dimension greater than 3 or 4.
An alternative approach proposed by Li et al. [22] is to allow the vector of dictionary elements $\Psi$ to be represented by a neural network, parameterized by weights $\theta_{\Psi}$; that is, $\Psi(\mathbf{x})=\Psi(\mathbf{x};\theta_{\Psi})$. This allows the basis to be optimized jointly along with the matrix approximation of the Koopman operator, $\mathbf{K}_{\tau}$, so that the objective in Eq. 18 becomes
@@ III.3 EDMD with dictionary learning
This approach is known as EDMD with dictionary learning (EDMD-DL). The original paper of Li et al [22] proposed an iterative procedure for solving Eq. 21 which alternates between two steps in which (1) the dictionary is fixed and the Koopman approximation $\mathbf{K}_{\tau}$ is optimized through the pseudoinverse calculation in Eq. 20 and then (2) the matrix $\mathbf{K}_{\tau}$ is fixed and the dictionary is optimized through stochastic gradient descent. Recently, Constante-Amores et al. showed that these steps can be combined by performing the least-squares calculation for $\mathbf{K}_{\tau}$ within the dictionary optimization step and backpropagating through this pseudoinverse-based solution of the least squares problem in optimizing the dictionary [23]. Other works have optimized the elements of $\mathbf{K}_{\tau}$ directly along with the dictionary by stochastic gradient descent, avoiding the least squares calculation entirely (e.g., [25, 24, 60]).
III.4 Forecasting with EDMD models
Here we discuss different variations of EDMD-DL for predicting the time-evolution of the states. First, we discuss the development of a projection operation which maps from the feature vector $\Psi$ to the states. Then we discuss how such a projection may be implemented within time-series prediction. This can either be done as a post-processing step of a full trajectory predicted linearly in the space of observables or by projecting back to the state space on each timestep, thereby formulating a nonlinear map on the state space.
III.4.1 Recovering $\mathbf{x}$ from $\Psi$
Once the dictionary $\Psi$ and the matrix operator $\mathbf{K}_{\tau}$ are determined, the time evolution of an observable $g_{\mathbf{c}}$ in $\mathcal{O}_{\mathbb{D}}$ can be predicted by expressing it in terms of the dictionary $\mathbb{D}$ as in Eq. 14 and then stepping the projection coefficients $\mathbf{c}$ forward in time with $\mathbf{K}_{\tau}$ as in Eq. 15. However, we are often more interested in forecasting the evolution of $\mathbf{x}$ along a single trajectory from an initial condition $\mathbf{x}_{0}$ at $t=0$. This can be done by using $\mathbf{K}_{t}$ to evolve the dictionary elements $\Psi(\mathbf{x})$ forward in time, as in the objective in Eq. 18,
@@ III.4.1 Recovering $\mathbf{x}$ from $\Psi$
Again, note that this will only be exact if the span of $\mathcal{O}_{\mathbb{D}}$ is Koopman invariant; otherwise, the residual can lead to significant error buildup in this prediction. While this method propagates the features $\Psi$ linearly, it is still necessary to recover the states $\mathbf{x}$ from the predictions of $\Psi$. That is, we need a mapping $\mathcal{P}:\mathbb{R}^{n}\to\mathbb{R}^{d}$, which effectively inverts the lifting $\Psi$, projecting the features back to the state space; i.e. $\mathbf{x}=\mathcal{P}\Psi(\mathbf{x})$. One common approach to this is to structure the dictionary so that the states $\mathbf{x}$ are included explicitly [22, 23]; that is,
@@
where $\tilde{\Psi}$ are dictionary elements which are nonlinear functions of $\mathbf{x}$. With this structure, the projection back to the state space takes the simple, linear form $\mathcal{P}\Psi(\mathbf{x})=\mathbf{P}\Psi(\mathbf{x})$ with $\mathbf{P}=\begin{bmatrix}\mathbf{I}_{n\times n}&\mathbf{0}_{n\times(k-n)}\end{bmatrix}$. Alternatively, if the states are not explicitly included in the dictionary, the mapping $\mathbf{P}$ can be learned as a linear mapping in a similar manner to $\mathbf{K}_{\tau}$; that is, by a least squares calculation for standard EDMD [61] or jointly by gradient descent for EDMD-DL. A more general approach is to assume that $\mathcal{P}$ is a nonlinear mapping and represent it as a decoder neural network, as in many Koopman autoencoder formulations [24, 25]. Otto and Rowley [25] performed comparisons of nonlinear reconstruction versus linear reconstruction of the observables from the features, showing that for some examples performance actually degrades by allowing for nonlinear reconstruction. For this reason, in our implementations of EDMD-DL, we will include the state explicitly in the dictionary, as in Eq. 23. As noted by Li, et al [22], including the states $\mathbf{x}$ as nontrainable dictionary elements also prevents the training procedure from learning the trivial solution, $\Psi(\mathbf{x})=\mathbf{0}$, which is a minimizer of Eq. 21. As pointed out by Otto and Rowley [25], it is also useful to normalize the error in the state prediction and the dictionary prediction in the loss function
$J(\mathbf{K}_{\tau};\theta_{\Psi})=\sum_{i=1}^{m}\left[\frac{\|\mathbf{y}_{i}-\mathbf{P}\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x}_{i};\theta_{\Psi})\|_{2}}{\|\mathbf{y}_{i}\|_{2}+\epsilon}+\lambda\,\frac{\|\tilde{\Psi}(\mathbf{y}_{i};\theta_{\Psi})-\tilde{\mathbf{P}}\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x}_{i};\theta_{\Psi})\|_{2}}{\|\tilde{\Psi}(\mathbf{y}_{i};\theta_{\Psi})\|_{2}+\epsilon}\right]$ (24)
@@
III.4.2 EDMD-DL with projection
@@
where the subscripts on $\Psi$ refer to values at timesteps evenly spaced by time interval, $\tau$. Iterating this for multiple timesteps allows us to obtain a trajectory in the feature space. Once this trajectory is obtained, the corresponding trajectory of the states, $\mathbf{x}$ can be obtained by mapping back to the state space using the operator $\mathcal{P}$. A key characteristic of Koopman-based approaches is that, as long as the operator $\mathcal{P}$ is linear, these methods allow for state trajectories to be predicted from an initial condition using only linear operations (following the initial nonlinear mapping to the feature space). This is key, as preserving linearity is necessary in order to apply linear methods for control [61].
@@
To understand why this subtle change offers such improved performance, we first note that Eq. 26 can be thought of as a nonlinear discrete-time model on the state space with a particular form, as depicted in Fig. 1. That is, the model in Eq. 26 is of the same form as Eq. 9 where
@@
So, in EDMD-DL, we learn a set of features $\Psi$ which is typically of higher dimension than the original state representation, $\mathbf{x}$ and whose time evolution is approximately linear. However, when we implement this model for state prediction with a reprojection on each step, we have essentially defined a nonlinear discrete time map on the state space of a form that first expands out to a high dimensional feature vector, followed by two linear mappings: the first, $\mathbf{K}_{\tau}^{\mathrm{T}}$, representing the time evolution and the second, $\mathbf{P}$, representing a projection back to the original state space. Since it is common for the output layer of a neural network to be linear (with no activation function following it), this model is essentially a neural network representation of the flow map in a fairly standard form. This observation is the heart of the present work.
Van Goor, et al. [48] recently examined this method as well, showing that there is a geometric interpretation of this step as a projection onto a manifold. To see this, we can think of the nonlinear feature map $\Psi$ as a coordinate mapping to a manifold in a higher dimensional space, $\Psi:\mathbb{R}^{n}\to\mathcal{M}_{\Psi}\subset\mathbb{R}^{k}$. That is, $\mathcal{M}_{\Psi}\subset\mathbb{R}^{k}$ is defined as the image of $\Psi$
$\mathcal{M}_{\Psi}=\left\{\,{\Psi}(\mathbf{x})\mid\mathbf{x}\in\mathbb{R}^{n}\,\right\}\subset\mathbb{R}^{k}\,.$ (28)
Then a trajectory in the state space, $\mathbf{x}(t)$, also defines a trajectory $\Psi(\mathbf{x}(t))$ on $\mathcal{M}_{\Psi}$. If the function space spanned by the elements of $\mathbb{D}$ is Koopman invariant, then the dynamics on $\mathcal{M}_{\Psi}$ can be represented exactly by a linear mapping
$\Psi(\mathbf{F}_{\tau}(\mathbf{x}))=\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x})\,.$ (29)
However, if $\mathcal{O}_{\mathbb{D}}$ is not Koopman invariant, the true operator $\mathcal{K}_{t}$ will map elements of $\mathcal{O}_{\mathbb{D}}$ out of this space, so a linear approximation of the dynamics on $\mathcal{M}_{\Psi}$ will not be exact. That is, the EDMD approximation $\mathbf{K}_{\tau}$ will map elements of $\mathcal{M}_{\Psi}$ off of this manifold even though $\mathcal{M}_{\Psi}$ is invariant under the dynamics by definition (i.e., $\Psi(\mathbf{F}_{\tau}(\mathbf{x}))\in\mathcal{M}_{\Psi}$). Performing a projection step in the forward rollout can be viewed as a projection back onto the manifold $\mathcal{M}_{\Psi}$ of the prediction using the EDMD approximation $\mathbf{K}_{\tau}^{\mathrm{T}}$ when the space $\mathcal{O}_{\mathbb{D}}$ is not Koopman invariant. By mapping back to the state space on each step and then re-lifting on the next step, it is ensured that the prediction remains on $\mathcal{M}_{\Psi}$. Specifically, $\Psi(\mathbf{P}\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x}))\in\mathcal{M}_{\Psi}$ even when $\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x})\notin\mathcal{M}_{\Psi}$.
III.5 Continuous-time models from EDMD

### v2

@@ III Numerical Methods
where $\mathbf{x}_{i}\equiv\mathbf{x}(t_{i})$, $\tau=t_{i+1}-t_{i}$ is the sampling interval, and $m$ is the number of data pairs. For the continuous time formulations considered below, it is not necessary to have a fixed time interval between datapoints in the dataset, but we will take it to be fixed here for consistency between methods. For each of the models considered, we will optimize the appropriate loss over the given dataset. Therefore, we assume that the dataset is sufficiently large and sampled from a measure on the state space (for example by collecting long trajectories, which effectively samples from a natural measure), so that useful approximations of the dynamics can be inferred from the data. Under this sampling condition and in the limit of large data, it can be shown that EDMD approximations converge to a Galerkin projection of the Koopman operator [14, 34, 35].
While we will be working with a clean, noise-free dataset of full-state observations uniformly spaced in time, we note that other works have considered extensions of the neural ODE framework to more realistic data scenarios. For example, in cases where data is gathered with irregular sampling times, it is still straightforward to train a neural ODE model [1], as the numerical integration can return model predictions at arbitrary points in time. Additionally, Ref. 36 showed that these sorts of neural ODE and discrete-time neural network models can be extended to cases where only partial state observations of a chaotic system are available by making use of Takens’ embedding theorem and time-delay embeddings. Other methods such as latent ODEs [1] and neural controlled differential equations [37] have been developed to extend neural ODE models to realistic scenarios, such as missing data, partial observations, irregular sampling, and incoming data. With regards to sensor noise, neural ODEs have been shown to perform well in cases of moderate noise added to the data without altering the training procedure [1, 38]. In cases of severe sensor noise, other steps could be taken, such as smoothing the data with a filter as a pre-processing step or adding additional regularization to the neural ODE training to discourage overfitting (see, for example, 38, 39, 40)
We will additionally be interested in fitting models using neural networks as function approximators, and those which preserve the Markovian structure assumed in the dynamical systems formulation above – that future time evolution is determined only from the present states. This requirement excludes several popular deep learning based time series modeling approaches, such as reservoir computing, LSTMs, and transformers.
Continuous-time and discrete-time models each have their own advantages and disadvantages. In many cases, especially when dealing with physical systems governed by continuous-time equations of motion, a continuous-time model is preferred as it more naturally represents continuous-time physics. Additionally, continuous-time models offer greater versatility in terms of temporal spacing of the training data and predictions, whereas discrete-time models are typically constrained to a single fixed timestep. It has also been shown that neural ODE methods tend to outperform discrete-time models in cases where the data is widely spaced in time [41], which is the case in many real-world data scenarios. However, in applications where model runtime is critical and predictions are needed only at a constant sampling frequency, discrete-time models can be advantageous. These models often provide faster inference since they generally require fewer function evaluations per timestep than neural ODEs, which typically use more sophisticated timestepping schemes.
@@ III.1 Neural ODEs
Developing a continuous-time model requires an approximation of the ODE vector field $\mathbf{f}(\mathbf{x})$ in Eq. 1. When $\mathbf{f}$ is approximated by a neural network $\mathbf{f}(\mathbf{x};\theta_{\mathbf{f}})$, the resulting model is called a neural ODE [1]. The parameters of a neural ODE are optimized to minimize the objective
@@
In practice, Eq. 10 is evaluated by numerically integrating from a given initial condition. In this work, we use a standard Dormand-Prince 4(5) algorithm implemented in the torchdiffeq [42] library using Python and PyTorch [43]. The objective in Eq. 9 is typically optimized by stochastic gradient descent, where the gradients with respect to the parameters $\theta_{\mathbf{f}}$ are determined by either solving an adjoint equation backward in time from $t_{i}+\tau$ to $t_{i}$ to obtain these sensitivities or by using modern automatic differentiation tools to backpropagate these gradients through the operations of a numerical ODE solver [1]. Once the neural ODE model is trained, it can be implemented to predict the time evolution of a system by simply integrating along the learned vector field using a numerical ODE solver from a given initial condition.
III.2 EDMD with dictionary learning
Extended dynamic mode decomposition (EDMD) [14] constructs a matrix approximation of the Koopman operator, $\mathbf{K}_{\tau}$ by solving a least squares problem to minimize the objective
@@ III.2 EDMD with dictionary learning
over the dataset of interest for a given vector of dictionary functions, $\Psi$. For full details on the EDMD minimization procedure, we refer the reader to Ref. 14.
A challenge of the EDMD approach is finding an appropriate dictionary on which to compute the operator $\mathbf{K}_{t}$, as the as the accuracy and convergence of the approximation depend heavily on this choice. Ideally, the subspace spanned by the dictionary should be Koopman invariant so that the residual can be driven to zero. To meet this requirement, it is necessary that the dictionary elements all lie in a space spanned by Koopman eigenfunctions [6, 10]. Additionally, we would like a dictionary from which the states can be recovered easily, and preferably linearly, for use in prediction applications. However, finding a dictionary that satisfies these properties is nontrivial, as the appropriate dictionary depends heavily on the underlying dynamical system and many common choices of dictionary do not scale well with the state dimension, making them challenging to apply to systems of dimension greater than 3 or 4.
An approach proposed by Li et al. [2] is to allow the vector of dictionary elements $\Psi$ to be represented by a neural network, parameterized by weights $\theta_{\Psi}$; that is, $\Psi(\mathbf{x})=\Psi(\mathbf{x};\theta_{\Psi})$. This allows the basis to be optimized jointly along with the matrix approximation of the Koopman operator, $\mathbf{K}_{\tau}$, so that the objective in Eq. 11 becomes
@@
This approach is known as EDMD with dictionary learning (EDMD-DL). The original paper of Li et al [2] proposed an iterative procedure for solving Eq. 12 which alternates between two steps in which (1) the dictionary is fixed and the Koopman approximation $\mathbf{K}_{\tau}$ is optimized through the EDMD and then (2) the matrix $\mathbf{K}_{\tau}$ is fixed and the dictionary is optimized through stochastic gradient descent. Recently, Constante-Amores et al. [16] showed that these steps can be combined by performing the least-squares calculation for $\mathbf{K}_{\tau}$ within the dictionary optimization step and backpropagating through both the lifting and the least-squares EDMD optimization of $\mathbf{K}_{\tau}$. Other works have optimized the elements of $\mathbf{K}_{\tau}$ directly along with the dictionary by stochastic gradient descent, avoiding the least squares calculation entirely (e.g., 18, 17, 44).
III.3 Forecasting with EDMD models
To forecast the evolution of the states, $\mathbf{x}$ from an initial condition, $\mathbf{x}_{0}$ at $t=0$, we begin by using $\mathbf{K}_{t}$ to evolve the dictionary elements $\Psi(\mathbf{x})$ forward in time, as in the objective in Eq. 11,
@@ III.3 Forecasting with EDMD models
While this method propagates the features $\Psi$ linearly, it is still necessary to recover the states $\mathbf{x}$ from the predictions of $\Psi$. That is, we need a mapping $\mathcal{P}:\mathbb{R}^{n}\to\mathbb{R}^{d}$, which effectively inverts the lifting $\Psi$, projecting the features back to the state space; i.e. $\mathbf{x}=\mathcal{P}\Psi(\mathbf{x})$. One common approach to this is to structure the dictionary so that the states $\mathbf{x}$ are included explicitly [2, 16]; that is,
@@
where $\tilde{\Psi}$ are dictionary elements which are nonlinear functions of $\mathbf{x}$. With this structure, the projection back to the state space takes the simple, linear form $\mathcal{P}\Psi(\mathbf{x})=\mathbf{P}\Psi(\mathbf{x})$ with $\mathbf{P}=\begin{bmatrix}\mathbf{I}_{n\times n}&\mathbf{0}_{n\times(k-n)}\end{bmatrix}$. Alternatively, if the states are not explicitly included in the dictionary, the mapping $\mathbf{P}$ can be learned as a linear mapping in a similar manner to $\mathbf{K}_{\tau}$; that is, by a least squares calculation for standard EDMD [45] or jointly by gradient descent for EDMD-DL. A more general approach is to assume that $\mathcal{P}$ is a nonlinear mapping and represent it as a decoder neural network, as in many Koopman autoencoder formulations [17, 18]. Otto and Rowley [18] performed comparisons of nonlinear reconstruction versus linear reconstruction of the observables from the features, showing that for some examples performance actually degrades by allowing for nonlinear reconstruction. For this reason, in our implementations of EDMD-DL, we will include the state explicitly in the dictionary, as in Eq. 14. Including the states $\mathbf{x}$ as nontrainable dictionary elements also prevents the training procedure from learning the trivial solution, $\Psi(\mathbf{x})=\mathbf{0}$, which is a minimizer of Eq. 12[2]. It is also useful to normalize the error in the state prediction and the dictionary prediction in the loss function[18]
$\begin{split}J(\mathbf{K}_{\tau};\theta_{\Psi})&=\sum_{i=1}^{m}\Bigg{[}\frac{\|\mathbf{y}_{i}-\mathbf{P}\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x}_{i};\theta_{\Psi})\|_{2}}{\|\mathbf{y}_{i}\|_{2}+\epsilon}\\ &\qquad\qquad+\lambda\,\frac{\|\tilde{\Psi}(\mathbf{y}_{i};\theta_{\Psi})-\tilde{\mathbf{P}}\mathbf{K}_{\tau}^{\mathrm{T}}\Psi(\mathbf{x}_{i};\theta_{\Psi})\|_{2}}{\|\tilde{\Psi}(\mathbf{y}_{i};\theta_{\Psi})\|_{2}+\epsilon}\Bigg{]}\end{split}$ (15)
@@
III.3.1 EDMD-DL with projection
@@
where the subscripts on $\Psi$ refer to values at timesteps evenly spaced by time interval, $\tau$. Iterating this for multiple timesteps allows us to obtain a trajectory in the feature space. Once this trajectory is obtained, the corresponding trajectory of the states, $\mathbf{x}$ can be obtained by mapping back to the state space using the operator $\mathcal{P}$. A key characteristic of Koopman-based approaches is that, as long as the operator $\mathcal{P}$ is linear, these methods allow for state trajectories to be predicted from an initial condition using only linear operations (following the initial nonlinear mapping to the feature space). This is key, as preserving linearity is necessary in order to apply linear methods for control, estimation, or other downstream tasks [45].
@@
To understand why this subtle change offers such improved performance, we first note that Eq. 17 can be thought of as a nonlinear discrete-time model on the state space with a particular form, as depicted in Fig. 1. That is, the model in Eq. 17 approximates the nonlinear discrete-time flow map on the state space as
@@
So, in EDMD-DL, a set of features $\Psi$ is learned, which is typically of higher dimension than the original state representation, $\mathbf{x}$, and whose time evolution is approximately linear. However, implementing this model for state prediction with a projection on each step defines a nonlinear discrete time map on the state space. This map first expands out to a high-dimensional feature vector, followed by two linear mappings: the first, $\mathbf{K}_{\tau}^{\mathrm{T}}$, represents the time evolution and the second, $\mathbf{P}$, represents a projection back to the original state space. Since it is common for the output layer of a neural network to be linear (with no activation function following it), this model is essentially a neural network representation of the flow map in a fairly standard form. This observation is the heart of the present work.
III.4 Continuous-time models from EDMD